5. Calculating tax incidence Suppose that the U.S. government decides to charge beer producers a tax. Before the tax, 25 billion cases of beer were sold every year at a price of $7 per case. After the tax, 19 billion cases of beer are sold every year; consumers pay $8 per case, and producers receive $4 per case (after paying the tax) The amount of the tax on a case of beer is$ per case. Of this amount, the burden that falls on consumers isS per case, and the burden that falls on producers is$per case. True or False: The effect of the tax on the quantity sold would have been smaller if the tax had been levied on consumers. True False
Answer #1

Amount of tax on a case of beer is = price that consumers pay-price the producers receive= 8-4=$4. Of this amount the burden of tax that fals on consumers=Price paid by consumers after tax-equilibrium price before tax=8-7=$1 and the burden that falls on producers=equilibrium price before tax-price received by producers after tax=7-4=$4

The effect tax would have been smaller if the tax would have been levied on consumers- FALSE, because it doesnt matter on whom the tax is levied, tax will affect the demand and supply in the same way.